OpenStudy (anonymous):

5√(x)+2=12 how do I go about finding x?

OpenStudy (anonymous):

\[12 = 5\sqrt{x} + 2\]\[10 = 5\sqrt{x}\]\[2 = \sqrt{x}\]\[2^2 = 4 = x\]

OpenStudy (anonymous):

5\[sqrt{x}\] +2 =12 5(x)^1/2=12-2 5x^1/2=10 x^1/2=10/5 x^1/2=2 square both sides x=2^2 x=4